OpenStudy (automaticloveletter):

Consider the following pair of equations: y = x + 4 y = −2x − 2 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

Mehek (mehek14):

ok first step you can substitute x + 4 into y for the second equation

Mehek (mehek14):

so instead of y in the second equation, you would have \(x+4=-2x-2\)

Mehek (mehek14):

now we can combine like terms add \(x+2x=\)? @automaticloveletter

Mehek (mehek14):

what do you get?

OpenStudy (automaticloveletter):

uhh I have no clue but 6? :/ I know i'm wrong

Mehek (mehek14):

no imagine x = \(1x\) what is \(1x+2x=1+2=?\)

OpenStudy (automaticloveletter):

3

Mehek (mehek14):

so \(3x\)

Mehek (mehek14):

now you have \(3x+4=-2\)

Mehek (mehek14):

now subtract \(-2-4=\)?

OpenStudy (automaticloveletter):

-6

Mehek (mehek14):

\(3x=-6\) divide \(\dfrac{-6}{3}\)

OpenStudy (automaticloveletter):

-2

Mehek (mehek14):

so x = -2

Mehek (mehek14):

now let's use the first equation and put -2 for x \(y=-2+4\\-2+4=?\)

OpenStudy (automaticloveletter):

2

Mehek (mehek14):

so y = 2 your answer is \((-2,2)\)

OpenStudy (automaticloveletter):

thank you for explaining

Mehek (mehek14):

your welcome \(\ddot\smile\)